OpenStudy (anonymous):

lim --> 3+ sec (pi(x)/6)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 3+} \sec \frac{ \pi(x) }{ 6 }\]

OpenStudy (anonymous):

you would get \[\sec(\frac{\pi}{2})\] which is undefined

OpenStudy (anonymous):

since you are going from the right, take a look at the graph of secant to see if you get infinity or minus infinity

OpenStudy (anonymous):

minus infinity

OpenStudy (anonymous):

How do you end up with sec (pi/2) though?

OpenStudy (anonymous):

\[\frac{3}{6}=\frac{1}{2}\]

OpenStudy (anonymous):

So once I have sec pi/2 , what happens?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

you are approaching it from the right, i.e. through numbers larger than \(\frac{\pi}{2}\)

OpenStudy (anonymous):

So I just put that the limit is minus infinity after that?